Find the solution by using the addition method. x+6y= -4, -x-4y= 2
Answer by stanbon(75887)

Find the solution by using the addition method.
x + 6y = -4
-x- 4y = 2
----------
Add and solve for "y":
2y = -2
y = -1
---
Solve for "x":
x + 6y = -4
x + 6*-1 = -4
x -6 = -4
x = 2
-------
Solution: (2,-1)
======================
Cheers,
Stan H.
======================